# The Rise of Ionized Gas Filaments in Early-Type Galaxies

### v1

@@ 5.2.1 HII Regions
These two distinct placements in the line diagnostic diagram suggest that the photoionization of filamentary nebulae, due to young massive stars, is more significant in low mass ETGs. While this sheds light on the mechanism responsible for powering their emission line spectra, this perhaps also helps distinguish which systems have an internal (cooling hot gas) vs. external (gas rich merger) origin of their warm filaments. In the latter process, the accretion of fresh, cold gas would birth young stars with radiation fields capable of ionizing surrounding gas. This process may be relevant in the low mass filamentary ETGs which sit in the leftmost region of Fig. 8. Whereas the high mass ETGs have their filaments formed from cooling hot gas. This is feasible for high mass systems since their steep gravitational potential allows them to retain a greater portion of their diffuse hot gas, relative to low mass ETGs.
@@ 5.2.2 AGN
In non star-forming galaxies, photoionization from AGN can produce line ratios similar to that of the diverse mechanisms powering LINERs. While AGN are certainly expected to photoionize gas restricted to the central regions (parsec scales), they cannot fully explain extended LINER emission. For example, Yan & Blanton (2012) found an [OIII]/[SII] ratio that increased with radius. They determine that this is explained by the ionization parameter, or the ratio between ionizing flux and gas density, also increasing with radius. This is physically motivated by the gas density having a shallower dependence on increasing radius, than that of the inverse-square law that radiation follows. An outward increasing ionization parameter is nonphysical if AGN are assumed to be the dominant ionization mechanism. Eracleous et al. (2010) similarly deduce that AGN are inadequate at producing LINER-like spectra, though from an energetics standpoint.
If AGN photoionization is the dominant source of powering the filamentary warm nebulae in our non-central ETGs, then we would expect their distributions in to significantly shift if we remove spaxels residing in the nucleur region. To test this, we only checked spaxels located at a distance greater than 0.3Re. In both the low and high mass filamentary sources, the peak of their extended spaxel distribution does not shift. Thus, AGN photons are likely not the main driver of the line emission in warm filaments.
@@ 5.2.3 Gas shocks
The role of gas shocks in powering the emission line spectra of ETGs cannot be ignored, given the wide range of phenomena that can produce them. Emission arising from collisional excitation within the heated downstream gas can result in LINER-like line ratios. If the shock is fast enough, the intense UV radiation field from the cooling downstream gas can photoionize the precursor gas leading to LINER/Seyfert-like spectra. Although, the relative importance of interstellar shocks in creating LINER-like spectra has long been debated. Shock models in Yan & Blanton (2012) produced line ratios that had velocity dependence opposite of that suggested by observations. Annibali et al. (2010) found their shock models to reproduce just the central emission line ratios, which could be explained by nuclear processes such as jet-driven outflows.
@@
We examine this relationship for all 19 of our filamentary ETGs in Fig. 9. To emphasize similar populations of spaxels, we apply K-means clustering to the data in each system. Each point in the figure represents the median of a K-means cluster in the -log([OI]/H) plane, and is color-coded by the median radial distance of the spaxels within that cluster. On average, these filamentary sources do not exhibit a tight correlation between velocity dispersion and [OI]/H, particularly in the more extended gas. While some systems show signatures of nuclear shocks (traced by centrally concentrated clusters with high velocity dispersion and line ratio values), it is likely not a dominant ionization mechanism across the bulk of the filamentary structures in our sample. Nevertheless, in the broader context of ETGs, we acknowledge that the role of gas shocks remains actively debated, as previous studies have demonstrated that shocks can reproduce LINER-like line ratios in extended gas (Lee et al., 2024).
@@
Photoionization from pAGB stars has long been thought to contribute to the emission line properties of ETGs (Trinchieri & di Serego Alighieri, 1991; Binette et al., 1994; Stasińska et al., 2008; Lagos et al., 2022). These are viable candidates since their harder radiation fields compared to younger stars are capable of reproducing line ratios observed in composite regions and LINERs. A strong correlation between emission line flux and stellar luminosity within emission line regions in ETGs (Macchetto et al., 1996; Sarzi et al., 2006, 2010) may be compelling evidence for this. Such a correlation would indicate that the production site of ionizing photons is distributed similarly to the stellar population. However, more recent studies using MaNGA observations of ETGs have challenged the feasibility of pAGB stars in photoionizing extended LINER gas, given discrepancies found between model predictions and their observed line ratios and temperature estimates (Yan, 2018; Lee et al., 2024).
For these reasons, we deduce that pAGB stars may have photoionized the warm filaments to some degree. However, given that the nebulae in these systems, unlike many ETGs with rotating gas disks, appear to be highly decoupled to the stellar population, an additional ionization source may be at play.
5.2.5 Self-irradiation cooling condensations
An interesting, often overlooked, source of optical line emission may be directly related to multiphase gas condensing out of the hot ISM. As thermally unstable gas cools from K down to K, X-ray and EUV photons will be emitted. Additionally, as described back in Sec. 5.1.2, lower temperature, higher density gas will condense out of the hot phase. The aforementioned high energy photons will irradiate the condensate, and can power its optical emission line spectrum. Studies of cooling flow clusters such as Voit & Donahue (1990) and Donahue & Voit (1991) have demonstrated how high energy photons from hot gas cooling could be reprocessed into H radiation, and produce their observed emission line spectra (see also Ferland et al. (2009); Polles et al. (2021)).
@@ 6 Conclusions
Using spatially resolved MUSE IFU observations of 126 non-central ETGs, we find that 54 host warm ionized gas nebulae. Out of these 54 sources, 35 have gas in the form of rotating disks. The remaining 19 display filamentary emission, similar to what is observed in massive group and cluster central galaxies. These MUSE observations along with Chandra X-ray observations allow us to infer that the dominant origin of these warm filaments is from the cooling, thermally-unstable hot halo in which they condensed out of. We also employ a set of thermodynamic and emission line analyses which provide insight on the physics within their ISM, thus allowing us to assess not only how filaments form, but what mechanisms regulate their existence and evolution. Our results can be summarized as:
@@
Our sample of 126 non-central ETGs, extracted from the 50MGC catalog, contains 54 that have warm ionized gas detected within their ISM. 35 host rotating warm gas disks, while 19 have filamentary warm gas. A portion of the observed filaments are compactly situated around their optical center, while others are significantly extended. These projected physical sizes of these filaments range from kpc.
@@
Unlike the non-central ETGs with rotating gas disks, nearly all (18/19) filamentary sources have nebulae that are misaligned with their stellar population. Stellar mass loss thus is likely not a dominant formation channel of the filaments. This observed misalignment may also suggest that external processes such as galaxy mergers and accretion may have contributed to a fraction of their warm ISM budget.
@@
All non-central ETGs with warm filaments (and with available Chandra observations) exhibit X-ray emission exceeding that expected from unresolved LMXBs. This suggests that they host an appreciable hot ISM. We posit that this hot atmosphere developed thermal instabilities allowing cooler, denser gas, such as the warm phase, to condense out of it. Previous studies of BCGs similarly deduce this to explain the formation of their filamentary nebulae. Our results of non-central ETGs, which reside in significantly less extreme and less dense environments, extend the formation mechanism of warm filaments to the general ETG population.
@@
NGC 4374, a non-central ETG within our sample has ample Chandra coverage, as well as pronounced hot filaments, which allow us to measure their electron pressure. We compare the electron pressure between the warm and hot filaments and find that the warm filaments has a higher thermal pressure than that of the hot filaments. This is opposite as to what is found in BCGs, thus suggesting that while filaments similarly form between BCGs and non-central ETGs, the physical conditions of their ISMs may be significantly different. Thus, how the filaments are regulated over time may vary between the two systems. Additionally we find a median X-rayH filament surface brightness ratio higher than that found in BCGs.
@@
Our emission line analysis using the optical line diagnostic diagram from Ji & Yan (2020) suggests that different mechanisms are controlling the line emission in non-central ETGs with rotating gas disks and those with filaments. Standard ionization sources such as AGN and gas shocks appear to be unable to describe the line emission in filamentary sources. Photoionization from pAGB stars may be capable of powering the filaments, but a lack of spatial and kinematic overlap between the stars and extended filaments may suggest that this is not the sole mechanism. Given that we expect there to be significant hot gas cooling in our filamentary sources, we posit that the high energy EUV photons and X-rays emanating from the cooling hot phase, may be radiating their warm condensate, and become reprocessed into optical emission. This advocates for future studies to focus on the role of cooling and self-irradiation in explaining the long-standing puzzle of the observed “composite” and LINER-like emission in ETGs.
@@ Acknowledgments
The authors thank the anonymous reviewer for their helpful comments on the manuscript. R.E., V.O., and Y.S. acknowledge support by NSF grant 2107711, NRAO grant SOSPA9-006, Chandra X-ray Observatory grants GO1-22126X, GO2-23120X, G01-22104X, NASA grants 80NSSC21K0714 and 80NSSC22K0856. V.O. acknowledges support from the DYCIT ESO-Chile Comite Mixto PS 1757, Fondecyt Regular 1251702, and DICYT PS 541.

### v2

@@ 5.2.1 HII Regions
These two distinct placements in the line-diagnostic diagram suggest that the photoionization of filamentary nebulae, due to young massive stars, is more significant in low mass ETGs. While this sheds light on the mechanism responsible for powering their emission line spectra, this perhaps also helps distinguish which systems have an internal (cooling hot gas) vs. external (gas rich accretion/merger) origin of their warm filaments. In the latter process, the accretion of fresh, cold gas would fuel star formation and provide radiation fields capable of ionizing surrounding gas. This process may be relevant in the low mass filamentary ETGs which sit in the leftmost region of Fig. 9. Consistent with this picture, ESO 375-041, the only ETG with warm gas filaments whose soft X-ray emission is not clearly above the unresolved-LMXB expectation, is also a low-mass () system. This suggests that at least some low-mass systems may host filaments without requiring a substantial hot-gas halo. Conversely, the warm gas filaments in the high-mass ETGs are more plausibly linked to a cooling origin, as this process is more viable in deeper gravitational potentials that are able to retain diffuse hot gas.
@@ 5.2.2 AGN
In non star-forming galaxies, photoionization from AGN can produce line ratios similar to that of the diverse mechanisms powering LINERs. While AGN are certainly expected to photoionize gas restricted to the central regions (parsec scales), they cannot fully explain extended LINER emission. For example, Yan and Blanton (2012) found an [OIII]/[SII] ratio that increased with radius. They determine that this is explained by the ionization parameter, or the ratio between ionizing flux and gas density, also increasing with radius. This is physically motivated by the gas density having a shallower dependence on increasing radius, than that of the inverse-square law that radiation follows. An outward increasing ionization parameter is nonphysical if AGN are assumed to be the dominant ionization mechanism. Eracleous et al. (2010) similarly deduce that AGN are inadequate at producing LINER-like spectra, though from an energetics standpoint.
If AGN photoionization is the dominant source of powering the filamentary warm nebulae in our non-central ETGs, then we would expect their distributions in to significantly shift if we remove spaxels residing in the nuclear region. To test this, we only checked spaxels located at a distance greater than 0.3Re. In both the low and high mass filamentary sources, the peak of their extended spaxel distribution does not shift. Thus, AGN photons are likely not the main driver of the line emission in warm filaments.
@@ 5.2.3 Gas shocks
The role of gas shocks in powering the emission line spectra of ETGs cannot be ignored, given the wide range of phenomena that can produce them. Emission arising from collisional excitation within the heated downstream gas can result in LINER-like line ratios. If the shock is fast enough, the intense UV radiation field from the cooling downstream gas can photoionize the precursor gas leading to LINER/Seyfert-like spectra. Although, the relative importance of interstellar shocks in creating LINER-like spectra has long been debated. Shock models in Yan and Blanton (2012) produced line ratios that had velocity dependence opposite of that suggested by observations. Annibali et al. (2010) found their shock models to reproduce just the central emission line ratios, which could be explained by nuclear processes such as jet-driven outflows.
@@
We examine this relationship for all 27 of our filamentary ETGs in Fig. 10. To emphasize similar populations of spaxels, we apply K-means clustering to the data in each system. Each point in the figure represents the median of a K-means cluster in the -log([OI]/H) plane, and is color-coded by the median radial distance of the spaxels within that cluster. On average, these filamentary sources do not exhibit a tight correlation between velocity dispersion and [OI]/H, particularly in the more extended gas. While some systems show signatures of nuclear shocks (traced by centrally concentrated clusters with high velocity dispersion and line ratio values), it is likely not a dominant ionization mechanism across the bulk of the filamentary structures in our sample. Nevertheless, in the broader context of ETGs, we acknowledge that the role of gas shocks remains actively debated, as previous studies have demonstrated that shocks can reproduce LINER-like line ratios in extended gas (Lee et al., 2024).
@@
Photoionization from pAGB stars has long been thought to contribute to the emission line properties of ETGs (Trinchieri and di Serego Alighieri, 1991; Binette et al., 1994; Stasińska et al., 2008; Lagos et al., 2022). These are viable candidates since their harder radiation fields compared to younger stars are capable of reproducing line ratios observed in composite regions and LINERs. A strong correlation between emission line flux and stellar luminosity within emission line regions in ETGs (Macchetto et al., 1996; Sarzi et al., 2006, 2010) may be compelling evidence for this. Such a correlation would indicate that the production site of ionizing photons is distributed similarly to the stellar population. However, more recent studies using MaNGA observations of ETGs have challenged the feasibility of pAGB stars in photoionizing extended LINER gas, given discrepancies found between model predictions and their observed line ratios and temperature estimates (Yan, 2018; Lee et al., 2024).
For these reasons, we deduce that pAGB stars may have photoionized the warm filaments to some degree. However, given that the nebulae in these systems, unlike many ETGs with rotating gas disks, tend to be highly decoupled to the stellar population, an additional ionization source may be at play.
5.2.5 Self-irradiated condensates
An interesting, often overlooked, source of optical line emission may be directly related to multiphase gas condensing out of the hot ISM. As thermally unstable gas cools from K down to K, X-ray and EUV photons will be emitted. Additionally, as described back in Sec. 5.1.2, lower temperature, higher density gas will condense out of the hot phase. The aforementioned high energy photons will irradiate the condensate, and can power its optical emission line spectrum. In this scenario, the hot gas and warm filaments should appear to be adjacent to each other. However, on average, our Chandra exposures are not deep enough to resolve the hot gas morphology at the scales of the warm filaments. Indeed, for the two filamentary ETGs in this sample with sufficiently deep Chandra observations, NGC 1266 and NGC 4374 (see Fig. 6), their hot gas appears to clearly overlap with their warm filaments (see Eskenasy et al., 2024 for more details on these two sources). Further evidence for spatial coincidence between the warm and hot gas phases in a subsample of non-cluster-central ETGs in the SAURON survey, including NGC 4374, is presented in Sarzi et al. (2010). This observational connection is consistent with a broader, multiphase cooling framework that has been explored in detail particularly in cooling-flow clusters. Voit and Donahue (1990) and Donahue and Voit (1991) have demonstrated how high energy photons from hot gas cooling could be reprocessed into H radiation, and produce the observed warm gas emission line spectra (see also Ferland et al., 2009; Sarzi et al., 2010; Polles et al., 2021).
@@ 6 Conclusions
Using spatially resolved MUSE IFU observations of 126 non-central ETGs, we find that 62 host warm ionized gas nebulae. Out of these 62 sources, 35 have gas in the form of rotating disks. The remaining 27 display filamentary emission, similar to what is observed in massive group and cluster central galaxies. These MUSE observations along with Chandra X-ray observations allow us to infer that cooling from a thermally-unstable hot halo is a primary channel for forming the warm filaments observed in non-central ETGs, while external acquisition may also contribute in some systems. We also employ a set of thermodynamic and emission line analyses which provide insight into the physics within their ISM, thus allowing us to assess not only how filaments form, but what mechanisms regulate their existence and evolution. Our results can be summarized as:
@@
Our sample of 126 non-central ETGs, extracted from the 50MGC catalog, contains 62 that have warm ionized gas detected within their ISM. Of these, 35 host rotating warm gas disks and 27 host filamentary warm gas. A portion of the observed filaments are compactly situated around their optical center, while others are significantly extended. Their projected physical sizes range from kpc.
@@
Unlike the majority of non-central ETGs with rotating gas disks, most (22/27) filamentary sources have nebulae that are misaligned with their stellar population. Stellar mass loss thus is likely not a dominant formation channel of the filaments. This misalignment may suggest that external processes such as galaxy mergers and accretion have contributed to a fraction of their warm-ISM budget.
@@
All filamentary ETGs with archival Chandra observations show soft X-ray emission that is consistent with exceeding (or, in one case, is indistinguishable from within the uncertainties) the expectation from unresolved LMXBs, indicating that systems with filamentary warm gas generally host an appreciable hot halo. We posit that thermal instabilities within this hot gas allow for cooler, denser phases, to condense out of it. This interpretation mirrors the leading model for filament formation in cool-core BCGs, and our results suggest that similar, cooling-driven pathways may be feasible in non-central ETGs that live in less extreme and less dense environments.
@@
NGC 4374, a non-central ETG within our sample has ample Chandra coverage, as well as pronounced hot filaments, which allow us to measure their electron pressure. We compare the electron pressure between the warm and hot filaments and find that the warm filaments have a higher thermal pressure than that of the hot filaments. This is opposite as to what is found in BCGs, thus suggesting that while filaments similarly form between BCGs and non-central ETGs, the physical conditions of their ISMs may be significantly different. Thus, how the filaments are regulated over time may vary between the two systems. Additionally, we find a median X-rayH filament surface brightness ratio higher than that found in BCGs.
@@
Our emission line analysis using the optical line-diagnostic framework of Ji and Yan (2020) indicates that different mechanisms are controlling the line emission in non-central ETGs with rotating gas disks and those with filaments. Standard, single-source explanations (such as AGN and gas shocks) appear to be unable to describe the line emission in the filaments. Photoionization from pAGB stars may be capable of powering the filaments, but a lack of spatial and kinematic overlap between the stars and extended filaments may suggest that this is not the sole mechanism. Given that we expect there to be significant hot gas in our filamentary sources, the high energy EUV photons and X-rays emanating from the cooling hot phase may be radiating their warm condensate, and, at least in part, power their emission lines. However, it is unclear if this is the dominant mechanism in all filamentary systems. This advocates for future studies to focus on the role of cooling and self-irradiation in explaining the long-standing puzzle of the observed “composite” and LINER-like emission in ETGs.
@@ Acknowledgments
The authors thank the referee for their thorough and helpful comments on the manuscript. R.E., V.O., and Y.S. acknowledge support by NSF grant 2107711, NRAO grant SOSPA9-006, Chandra X-ray Observatory grants GO1-22126X, GO2-23120X, G01-22104X, NASA grants 80NSSC21K0714 and 80NSSC22K0856. V.O. acknowledges support from the DYCIT ESO-Chile Comite Mixto PS 1757, Fondecyt Regular 1251702, and DICYT PS 541.